# Host Galaxy Demographics Of Individually Detectable Supermassive Black-hole Binaries with Pulsar Timing Arrays

### v1

@@ 3 PTA Detection Probability
Having tracked the formation of SMBHBs in Illustris, then post-processed their dynamical evolution through several stages of environmental interaction until the observed GW frequency falls within the frequency regime of PTA sensitivity, we are left with $N_{b}=3947$ binaries and host galaxies. We now describe how Monte Carlo realizations of SMBHB populations are synthesized from this sample, for which we mostly follow the scheme outlined in Kelley et al. [53]. We then discuss how we assess the probability of detecting GWs from individual binary systems in PTAs.
3.1 Generating SMBHB Populations
To synthesize Monte Carlo realizations of GW-emitting SMBHB populations, we consider the build-up of GW signals from individual systems with chirp mass $\mathcal{M}$ at different observer-frame frequencies $f$, such that the GW characteristic strain can be calculated as,
@@ 3.1 Generating SMBHB Populations
The discretization of subsection 3.1 over our finite number of sources from the Illustris simulation turns the integral over number density into a sum over sources within the Illustris volume ($V_{\mathrm{ill}}$):
@@
which corresponds to a sum over all binaries $i$ at all simulation time-steps $j$. The volume scaling factor $\Lambda_{ij}\equiv\Delta V_{ij}/V_{\mathrm{ill}}$ is the number of astrophysical SMBHBs in the observer’s past light cone that is represented by each simulated Illustris binary. This depends on the integration step-size $\Delta z_{ij}$,
@@
Our subsequent scheme now differs slightly from Kelley et al. [53], which associates each Illustris SMBHB with a weight drawn from a Poisson distribution $\mathcal{P}(\Lambda_{ij})$, then sums the squared characteristic strain to produce a particular realization of a GWB. Instead, we construct a kernel density estimator [^3] Using the kalepy package [51]. (KDE) with $\Lambda_{ij}$ as the weight of each simulated binary. This KDE is calculated across $9$ properties of the $N_{b}$ systems, including those of the SMBHB: total mass $M$, mass ratio, $q$, redshift $z$, GW frequency $f$, and those of the host galaxy: $g-r$ color, absolute magnitude ($M_{r}$), stellar metallicity ($Z$), specific star-formation rate (sSFR), and stellar mass ($M_{\star}$). Each of these quantities are then resampled from the KDE using reflective boundary conditions, either determined by physical reasons (e.g. mass ratio $0<q\leq 1$), or limited to the parameter bounds from Illustris (e.g. restricting total masses to be between the minimum and maximum resolved in the simulation volume). Thus our distributions of properties are conservative, as they don’t extrapolate beyond the limitations of the simulation.
Having constructed our KDE across the $9$-dimensional space of system properties, we re-sample $10^{3}$ realizations. We draw $\sim 2.5\times 10^{6}$ systems in each realization, corresponding to the sum of the $\Lambda_{ij}$ factors, i.e., the total number of astrophysical binaries expected in the observer’s past light cone. Using these random draws of system properties, we calculated the characteristic strain for each binary, and the GW background spectrum of each realization. In the left panel of Figure 2 we show the characteristic strain spectra for five of our $10^{3}$ realizations, contrasted with the idealized power-law behavior of $\propto f^{-2/3}$ [80]. As expected, GWB spectra from finite-number Monte Carlo realizations depart from the expected behavior at higher frequencies due to low occupancy factors [94]. On the right panel of Figure 2 we show just one realization, where we now exclude the loudest source in each frequency bin as a candidate individually-resolvable SMBHB. The remnant GWB signal from all other binaries is treated as a self-noise term in the determination of the loudest source’s GW-detection probability. We now discuss how this detection probability is calculated.
@@ 3.2 PTA Detection Statistics
Our scheme for assessing the PTA detection of GW signals from individual SMBHB systems uses the $\mathcal{F}_{e}$ statistic [33]. This statistic is based on the assumption that typical binary orbital evolution timescales are shorter than pulsar-Earth light-travel times, such that the frequencies of pulsar-terms will all be different and will lie below that of the common Earth-term frequency. Hence only the Earth-term waveform is used as a matched-filtering template when constructing the PTA likelihood. By forming a ratio of the likelihood in the presence of a signal to one in the absence of a signal, then analytically maximizing this over terms involving the initial GW phase $\Phi_{0}$, GW polarization angle $\psi$, binary inclination angle, $\iota$, and GW amplitude $A$, we can form a statistic that is only a function of GW sky-location $\{\theta,\phi\}$ and frequency $f$. In the absence of a signal, this statistic is distributed as [33]
@@
where $\bar{\mathcal{F}_{e}}$ is the threshold value of the statistic corresponding to the chosen FAP $\alpha_{0}$, and $N$ accounts for a number of independent trials of waveform templates that must be performed in order to determine $\{\theta,\phi,f\}$. Following Rosado & Sesana [87], we set $N=10^{4}$. We may then numerically solve Equation 6 to find the threshold value of $\bar{\mathcal{F}_{e}}$ for our chosen value of $\alpha_{0}=0.001$.
@@
where $\Phi_{T}=2\pi fT$ for a timespan of pulsar observations, $T$; $F_{j}^{+,\times}$ are the GW polarization response functions for each pulsar, which are a function of the binary’s and pulsar’s sky location; $a$ and $b$ are functions of the binary orbital inclination– see Rosado & Sesana [87] for full details and definitions. Finally, we calculate $S_{j}$, the noise power-spectral density of the timing residuals in each pulsar, given by
@@ 3.3 A Near-future PTA
Pulsar noise characteristics are selected to be currently realistic, albeit conservative as timing precision will improve in real datasets in the future. We adopt values from scenario-$(c)$ in Kelley et al. [56], which for all pulsars has $\sigma=0.3\mu\mathrm{s}$, $A_{\mathrm{red}}=10^{-15}$, and $\gamma_{\mathrm{red}}=4.5$. Combined, this corresponds to a total strain noise (i.e., not due to GWs) of $h_{c}^{N}=6.7\times 10^{-15}$ at $f=0.1$ yr-1. Obviously in real PTAs the noise characteristics are highly heterogeneous, with some pulsars having much better timing precision or lower-frequency sensitivity than others. Likewise, pulsar timing precision improves over time as new radio-timing hardware is brought online or as the duration of timing observations increases. Thus as a first test of our framework to weight SMBHB distributions by their detection probability, we also assess the impact of PTA noise characteristics.
Figure 4 shows DP-weighted distributions of binary chirp mass, corresponding to an average over our $10^{3}$ population realizations. Within each realization, the binary with the largest strain in each frequency resolution bin is selected, and its DP averaged over $10^{3}$ random draws of possible sky locations, initial phases, and orbital inclinations. This averaged DP is used as a weighting in our distributions. As expected, these favor high chirp-mass values, yet not the highest, since those systems at $\gtrsim 10^{10}M_{\odot}$ are exceedingly rare and evolve rapidly through the PTA band. We vary the pulsar noise characteristics across $4$ combinations of $\sigma=[0.1,0.3]~{}\mu\mathrm{s}$ and $A_{\mathrm{red}}=[0,10^{-15}]$. Surprisingly, we find that the white noise level has the largest overall impact on these distributions, where increased white noise pushes the most likely binary chirp masses to be higher, where they produce stronger, more detectable signals against the raised noise. White noise is broadband, whereas red noise is (by definition) lower frequency. Hence this also indicates that most detectable SMBHBs are found at higher frequencies in the PTA band where the red noise and GWB do not dominate sensitivity.
Figure 4: Detection-probability–weighted chirp mass distributions for SMBHBs are compared for varying PTA noise characteristics. As expected, higher noise across all frequencies leads to the distribution of detectable binary systems shifting to larger chirp masses which produce higher-amplitude signals. We adopt the $\sigma_{\mathrm{WN}}=300$ ns, $A_{\mathrm{RN}}=10^{-15}$ noise characteristics for the rest of this paper.
@@ 4 Results
Table 1: Median and $90\%$ range of the observable properties of host galaxies and constituent SMBHBs for GW-detectable systems. Values are derived from our GW detection-probability–weighted distributions.
In this section we compare the population of host galaxies weighted by the GW detection-probability of their constituent SMBHBs, against the population of all binary host galaxies. We remark on how different galaxy observables for hosts of detectable binaries are different from the population of all binary hosts, as well as all galaxies in general. Table 1 summarizes our key findings for the median and $90\%$ range of GW detection-weighted population distributions for host galaxies and constituent SMBHBs.
Figure 5: Distribution of SMBHB properties for all binaries in our Illustris selection of systems emitting GWs in the PTA band. The blue dashed line corresponds to the unweighted population properties, while solid black lines correspond to weighting systems by their GW detection probability. The left panel shows total binary mass, the center panel shows the redshift, and the right panel shows binary mass ratio, $M_{2}/M_{1}$. Note the different vertical scales in each panel for ease of viewing.
@@
Figure 7: Distribution of host-galaxy properties for all binaries in our Illustris selection of systems emitting GWs in the PTA band. The blue dashed line corresponds to the unweighted host-galaxy population properties, while solid black lines correspond to weighting the host-galaxy distributions by the detectability of their constituent SMBHB. The left panel shows stellar mass, the center panel shows stellar metallicity with respect to the Sun, and the right panel shows specific star formation rate.
4.1 SMBHB Properties
Figure 5 shows the comparison between the distribution of SMBHB properties for all binaries emitting GWs in the PTA band to those of GW-detectable systems. In the left panel, the unweighted population is dominated by lower total-masses, but the selection effects of GW detectability skews this distribution toward higher values around $\sim 10^{10}M_{\odot}$. Likewise, with GW signals having higher strain amplitudes for closer systems (in inverse proportion to their distance), the bulk of the distribution of detectable binaries is $z<1$, peaking at $z\sim 0.5$ as shown in the center panel. Finally, the right panel shows that the distribution of mass ratios for all binaries is skewed toward asymmetric systems. In Illustris, the overall distribution of mass-ratios is relatively flat in log-space. While the binary evolution model does lead to larger fractions of extreme mass-ratio systems stalling and not reaching the PTA frequency regime, a noticeable fraction do still reach these frequencies [52]. In contrast, the detection-weighted distribution peaks around $q\sim 0.4$. One can understand this by recalling that $h\propto\mathcal{M}_{\mathrm{chirp}}^{5/3}\propto qM_{\mathrm{total}}/(1+q)^{2}$, such that for fixed total mass the strain amplitude scales approximately linearly with $q$ at small-$q$ values. Hence GW detectability preferentially selects more equal mass binary systems. We finish this exploration of SMBHB properties with Figure 6, showing contrasting distributions of GW observables in the form of observer-frame chirp mass and observed GW frequency. The distributions are notably separated, demonstrating how the requirements of GW detectability selects for systems in the high observer-frame chirp mass tail, and preferring observed frequencies $\sim 5$ nHz.
@@ 4.2 Host Galaxy Properties
In Figure 7 we show the analog of Figure 5 for several properties of host galaxies. In the left panel we see that the peak of the stellar mass distribution shifts higher by almost two orders of magnitude from $\sim 10^{10}M_{\odot}$ to $\sim 10^{12}M_{\odot}$ when weighting host galaxies by GW detection probability. In the center panel we show galaxy stellar metallicity, which is a mass-weighted average metallicity of stars within twice the stellar half-mass radius of each Illustris subhalo. The most detectable SMBHBs are found in host galaxies where the stellar metallicity is approximately $2Z_{\odot}$. This is in notable contrast to the overall distribution of binary hosts, which peaks nearer to $0.6Z_{\odot}$. Galaxies with higher stellar metallicity tend to be older; metallicity increases with subsequent generations of stars, as additional metals are formed through nuclear burning, and ejected into the environment through supernovae and winds to then form the next generation of stars.
@@
Older galaxies are also more likely to have gone through major mergers and thus are more likely to host a SMBHB system that is massive enough to be detectable in GWs by PTAs. The fact that the distribution of stellar metallicity becomes noticeably shifted upon weighting by GW detection probability means that it could be one of the best properties for ranking a galaxy’s likelihood of hosting a GW-detectable SMBHB [109]. Finally, in the right panel of Figure 7 we show specific star formation rate, measuring the rate of star formation per unit stellar mass. Binary hosts in general peak around $\sim 10^{-9}\mathrm{yr}^{-1}$, which drops by approximately an order of magnitude or more for galaxies hosting GW-detectable binaries.
@@
Figure 8: Galaxy properties in the color-magnitude plane for absolute magnitude in $r$-band and $g-r$ color. As in Figure 5 and Figure 7, the blue dashed lines indicate the distribution for all binaries while the black solid lines correspond to the detection-weighed distribution. Orange-dotted lines represent the distribution for all galaxies. Contours enclose $68\%$ and $95\%$ of probability, respectively. The galaxies hosting the most GW-detectable SMBHBs tend to be redder (high in $g-r$ color) and brighter (more negative in absolute magnitude). The red dash-dot line indicates the decision boundary of a linear support vector machine classifier between the black-solid and orange-dotted distributions.
@@
We now turn to galaxy photometric observables, with Figure 8 showing the joint distributions of absolute magnitude in the $r$ band, $M_{r}$ (as an indicator of inherent galaxy brightness) versus $g-r$ galaxy color. As before, blue-dashed and black-solid represent the unweighted distribution for all binary hosts and the distribution weighted by the GW detectability of the constituent SMBHB, respectively. The orange-dotted line shows the distribution in the color-magnitude plane for all galaxies, not just binary hosts. Immediately obvious is that galaxies harboring SMBHBs that are detectable in GWs by PTAs tend to be brighter and redder than typical binary hosts and typical galaxies overall. This also implies that these galaxies are quenched ellipticals. We observe that the distribution of binary host properties overlaps with those of all galaxies and the hosts of GW-detectable binaries (as one would expect), but that the compounding requirements of a galaxy hosting a binary that is also detectable in GWs by PTAs leads to virtually negligible overlap in the black-solid and orange-dotted distributions. It is worth noting that the difference between unweighted binary-hosts and all galaxies is partially due to selection cuts: recall that in Illustris SMBH are seeded into halos above a certain mass, which will correlate with brightness.
@@
The red dash-dotted line in Figure 8 indicates the decision boundary for a linear support vector machine (SVM) classifier [e.g. 98] between these latter two distributions, achieving virtually perfect classification (i.e., the area under the receiver operator characteristic curve was $1.0$) on $20\%$ of sampled systems that were held out from training. [^4] We note that the SVM was trained on equal numbers of sampled systems from each distribution, such that the classification was based on the distribution of properties and not rates. By contrast, a similar SVM trained to discriminate between binary hosts and all galaxies achieves a true positive rate of $\sim 65\%$ with a false positive rate of $0.1\%$, due to the larger (visible) overlap in those distributions. A full systematic investigation of the optimal classification scheme for which galaxy observables can aid in discriminating the hosts of detectable SMBHBs is beyond the scope of this paper, but our proof-of-principle analysis here does demonstrate the feasibility of doing so with color–magnitude and other galaxy information.

### v2

@@ 3 PTA Detection Probability
Having tracked the formation of MBHBs in Illustris, then post-processed their dynamical evolution through several stages of environmental interaction until the observed GW frequency falls within the frequency regime of PTA sensitivity, we are left with $N_{b}=3947$ binaries and host galaxies. We now describe how Monte Carlo realizations of MBHB populations are synthesized from this sample, for which we mostly follow the scheme outlined in Kelley et al. [62]. We then discuss how we assess the probability of detecting GWs from individual binary systems in PTAs.
3.1 Generating MBHB Populations
To synthesize Monte Carlo realizations of GW-emitting MBHB populations, we consider the build-up of GW signals from individual systems with chirp mass $\mathcal{M}$ at different observer-frame frequencies $f$, such that the GW characteristic strain can be calculated as,
@@ 3.1 Generating MBHB Populations
The discretization of subsection 3.1 over our finite number of sources from the Illustris simulation turns the integral over number density into a sum over sources within the Illustris volume ($V_{\mathrm{ill}}=(106.5$ $\mathrm{Mpc})^{3}$):
@@
which corresponds to a sum over all binaries $i$ at all simulation time-steps $j$. The volume scaling factor $\Lambda_{ij}\equiv\Delta V_{ij}/V_{\mathrm{ill}}$ is the number of astrophysical MBHBs in the observer’s past light cone that is represented by each simulated Illustris binary. This depends on the integration step-size $\Delta z_{ij}$,
@@
Our subsequent scheme now differs slightly from Kelley et al. [62], which associates each Illustris MBHB with a weight drawn from a Poisson distribution $\mathcal{P}(\Lambda_{ij})$, then sums the squared characteristic strain to produce a particular realization of a GWB. Instead, we construct a kernel density estimator [^4] Using the kalepy package [59]. (KDE) with $\Lambda_{ij}$ as the weight of each simulated binary. This KDE is calculated across $9$ properties of the $N_{b}$ systems, including those of the MBHB: total mass $M$, mass ratio, $q$, redshift $z$, GW frequency $f$, and those of the host galaxy: $g-r$ color, absolute magnitude ($M_{r}$), stellar metallicity ($Z$), specific star-formation rate (sSFR), and stellar mass ($M_{\star}$). Each of these quantities are then resampled from the KDE using reflective boundary conditions, either determined by physical reasons (e.g. mass ratio $0<q\leq 1$), or limited to the parameter bounds from Illustris (e.g. restricting total masses to be between the minimum and maximum resolved in the simulation volume). Thus our distributions of properties are conservative, as they don’t extrapolate beyond the limitations of the simulation.
Having constructed our KDE across the $9$-dimensional space of system properties, we re-sample $10^{3}$ realizations. We draw $\sim 2.5\times 10^{6}$ systems in each realization, corresponding to the sum of the $\Lambda_{ij}$ factors, i.e., the total number of astrophysical binaries expected in the observer’s past light cone. Using these random draws of system properties, we calculated the characteristic strain for each binary, and the GW background spectrum of each realization. In the left panel of Figure 1 we show the characteristic strain spectra for five of our $10^{3}$ realizations, contrasted with the idealized power-law behavior of $\propto f^{-2/3}$ [92]. As expected, GWB spectra from finite-number Monte Carlo realizations depart from the expected behavior at higher frequencies due to low occupancy factors [107]. On the right panel of Figure 1 we show just one realization, where we now exclude the loudest source in each frequency bin as a candidate individually-resolvable MBHB. The remnant GWB signal from all other binaries is treated as a self-noise term in the determination of the loudest source’s GW-detection probability. We now discuss how this detection probability is calculated.
@@ 3.2 PTA Detection Statistics
Our scheme for assessing the PTA detection of GW signals from individual MBHB systems uses the $\mathcal{F}_{e}$ statistic [39]. This statistic is based on the assumption that typical binary orbital evolution timescales are shorter than pulsar-Earth light-travel times, such that the frequencies of pulsar-terms will all be different and will lie below that of the common Earth-term frequency. Hence only the Earth-term waveform is used as a matched-filtering template when constructing the PTA likelihood. By forming a ratio of the likelihood in the presence of a signal to one in the absence of a signal, then analytically maximizing this over terms involving the initial GW phase $\Phi_{0}$, GW polarization angle $\psi$, binary inclination angle, $\iota$, and GW amplitude $A$, we can form a statistic that is only a function of GW sky-location $\{\theta,\phi\}$ and frequency $f$. In the absence of a signal, this statistic is distributed as [39]
@@
where $\bar{\mathcal{F}_{e}}$ is the threshold value of the statistic corresponding to the chosen FAP $\alpha_{0}$, and $N$ accounts for a number of independent trials of waveform templates that must be performed in order to determine $\{\theta,\phi,f\}$. Following Rosado & Sesana [99], we set $N=10^{4}$. We may then numerically solve Equation 6 to find the threshold value of $\bar{\mathcal{F}_{e}}$ for our chosen value of $\alpha_{0}=0.001$, which was previously adopted by Rosado & Sesana [99] and corresponds to $\sim 3$ $\sigma$ significance. While our results would vary quantitatively with different FAP, we expect that our conclusions would not qualitatively change. Detectable gravitational-wave signals from SMBHBs favor massive binaries, high frequencies, and small distances. However, these signal selection effects are convolved with the underlying rate distribution of massive black holes in the Universe. Lowering the FAP would likely select for stronger signals and more extreme system properties, but our choice of FAP here is a reasonable one to avoid spurious noise sources.
@@
where $\Phi_{T}=2\pi fT$ for a timespan of pulsar observations, $T$; $F_{j}^{+,\times}$ are the GW polarization response functions for each pulsar, which are a function of the binary’s and pulsar’s sky location; $a$ and $b$ are functions of the binary orbital inclination– see Rosado & Sesana [99] for full details and definitions. We distribute all binary positions to be statistically isotropic on the sky, which is well justified given that the influence of galaxy clustering on small scales is likely to be a minor consideration for PTA GW searches [4]. Finally, we calculate $S_{j}$, the noise power-spectral density of the timing residuals in each pulsar, given by
@@ 3.3 A Near-future PTA
Pulsar noise characteristics are selected to be currently realistic, albeit conservative as timing precision will improve in real datasets in the future. We adopt values from scenario-$(c)$ in Kelley et al. [66], which for all pulsars has $\sigma=0.3\mu\mathrm{s}$, $A_{\mathrm{red}}=10^{-15}$, and $\gamma_{\mathrm{red}}=4.5$. Combined, this corresponds to a total strain noise (i.e., not due to GWs) of $h_{c}^{N}=6.7\times 10^{-15}$ at $f=0.1$ yr-1. Obviously in real PTAs the noise characteristics are highly heterogeneous, with some pulsars having much better timing precision or lower-frequency sensitivity than others. Likewise, pulsar timing precision improves over time as new radio-timing hardware is brought online or as the duration of timing observations increases. Thus as a first test of our framework to weight MBHB distributions by their detection probability, we also assess the impact of PTA noise characteristics.
Figure 3 shows DP-weighted distributions of binary chirp mass, corresponding to an average over our $10^{3}$ population realizations. Within each realization, the binary with the largest strain in each frequency resolution bin is selected, and its DP averaged over $10^{3}$ random draws of possible sky locations, initial phases, and orbital inclinations. This averaged DP is used as a weighting in our distributions. As expected, these favor high chirp-mass values, yet not the highest, since those systems at $\gtrsim 10^{10}M_{\odot}$ are exceedingly rare and evolve rapidly through the PTA band. We vary the pulsar noise characteristics across $4$ combinations of $\sigma=[0.1,0.3]~{}\mu\mathrm{s}$ and $A_{\mathrm{red}}=[0,10^{-15}]$. Surprisingly, we find that the white noise level has the largest overall impact on these distributions, where increased white noise pushes the most likely binary chirp masses to be higher, where they produce stronger, more detectable signals against the raised noise. White noise is broadband, whereas red noise is (by definition) lower frequency. Hence this also indicates that most detectable MBHBs are not found at the very lowest frequencies, but at slightly higher frequencies in the PTA band where the red noise and GWB do not dominate sensitivity.
Figure 3: Detection-probability–weighted chirp mass distributions for MBHBs are compared for varying PTA noise characteristics. As expected, higher noise across all frequencies leads to the distribution of detectable binary systems shifting to larger chirp masses which produce higher-amplitude signals. We adopt the $\sigma_{\mathrm{WN}}=300$ ns, $A_{\mathrm{RN}}=10^{-15}$ noise characteristics for the rest of this paper.
@@ 4 Results
Table 1: Median and $90\%$ range of the observable properties of host galaxies and constituent MBHBs for GW-detectable systems. Values are derived from our GW detection-probability–weighted distributions.
In this section we compare the population of host galaxies weighted by the GW detection-probability of their constituent MBHBs, against the population of all binary host galaxies. We remark on how different galaxy observables for hosts of detectable binaries are different from the population of all binary hosts, as well as all galaxies in general. Table 1 summarizes our key findings for the median and $90\%$ range of GW detection-weighted population distributions for host galaxies and constituent MBHBs.
Figure 4: Distribution of MBHB properties for all binaries in our Illustris selection of systems emitting GWs in the PTA band. The blue dashed line corresponds to the unweighted population properties, while solid black lines correspond to weighting systems by their GW detection probability. The left panel shows total binary mass, the center panel shows the redshift, and the right panel shows binary mass ratio, $M_{2}/M_{1}$. Note the different vertical scales in each panel for ease of viewing.
@@
Figure 6: Distribution of host-galaxy properties for all binaries in our Illustris selection of systems emitting GWs in the PTA band. The blue dashed line corresponds to the unweighted host-galaxy population properties, while solid black lines correspond to weighting the host-galaxy distributions by the detectability of their constituent MBHB. The left panel shows stellar mass, the center panel shows stellar metallicity with respect to the Sun, and the right panel shows specific star formation rate.
4.1 MBHB Properties
Figure 4 shows the comparison between the distribution of MBHB properties for all binaries emitting GWs in the PTA band to those of GW-detectable systems. In the left panel, the unweighted population is dominated by lower total-masses, but the selection effects of GW detectability skews this distribution toward higher values around $\sim 10^{10}M_{\odot}$. Likewise, with GW signals having higher strain amplitudes for closer systems (in inverse proportion to their distance), the bulk of the distribution of detectable binaries is $z<1$, peaking at $z\sim 0.5$ as shown in the center panel. [^5] Indeed, even at $z\sim 1$, ultra-massive binary BHs in massive clusters could be a source of resolvable signals [123]. Finally, the right panel shows that the distribution of mass ratios for all binaries is skewed toward asymmetric systems. In Illustris, the overall distribution of mass-ratios is relatively flat in log-space. While the binary evolution model does lead to larger fractions of extreme mass-ratio systems stalling and not reaching the PTA frequency regime, a noticeable fraction do still reach these frequencies [61]. In contrast, the detection-weighted distribution peaks around $q\sim 0.4$. One can understand this by recalling that $h\propto\mathcal{M}_{\mathrm{chirp}}^{5/3}\propto qM_{\mathrm{total}}/(1+q)^{2}$, such that for fixed total mass the strain amplitude scales approximately linearly with $q$ at small-$q$ values. Hence GW detectability preferentially selects more equal mass binary systems. We finish this exploration of MBHB properties with Figure 5, showing contrasting distributions of GW observables in the form of observer-frame chirp mass and observed GW frequency. The distributions are notably separated, demonstrating how the requirements of GW detectability selects for systems in the high observer-frame chirp mass tail, and preferring observed frequencies $\sim 5$ nHz.
@@ 4.2 Host Galaxy Properties
In Figure 6 we show the analog of Figure 4 for several properties of host galaxies. In the left panel we see that the peak of the stellar mass distribution shifts higher by almost two orders of magnitude from $\sim 10^{10}M_{\odot}$ to $\sim 10^{12}M_{\odot}$ when weighting host galaxies by GW detection probability. In the center panel we show galaxy stellar metallicity, which is a mass-weighted average metallicity of stars within twice the stellar half-mass radius of each Illustris subhalo. The most detectable MBHBs are found in host galaxies where the stellar metallicity is approximately $2Z_{\odot}$. This is in notable contrast to the overall distribution of binary hosts, which peaks nearer to $0.6Z_{\odot}$. Galaxies with higher stellar metallicity tend to be older; metallicity increases with subsequent generations of stars, as additional metals are formed through nuclear burning, and ejected into the environment through supernovae and winds to then form the next generation of stars.
@@
Older galaxies are also more likely to have gone through major mergers and thus are more likely to host a MBHB system that is massive enough to be detectable in GWs by PTAs. The fact that the distribution of stellar metallicity becomes noticeably shifted upon weighting by GW detection probability means that it could be one of the best properties for ranking a galaxy’s likelihood of hosting a GW-detectable MBHB [125]. Finally, in the right panel of Figure 6 we show specific star formation rate, measuring the rate of star formation per unit stellar mass. Binary hosts in general peak around $\sim 10^{-9}\mathrm{yr}^{-1}$, which drops by approximately an order of magnitude or more for galaxies hosting GW-detectable binaries.
@@
Figure 7: Galaxy properties in the color-magnitude plane for absolute magnitude in $r$-band and $g-r$ color. As in Figure 4 and Figure 6, the blue dashed lines indicate the distribution for all binaries while the black solid lines correspond to the detection-weighed distribution. Orange-dotted lines represent the distribution for all galaxies. Contours enclose $68\%$ and $95\%$ of probability, respectively. The galaxies hosting the most GW-detectable MBHBs tend to be redder (high in $g-r$ color) and brighter (more negative in absolute magnitude). The red dash-dot line indicates the decision boundary of a linear support vector machine classifier between the black-solid and orange-dotted distributions.
@@
We now turn to galaxy photometric observables, with Figure 7 showing the joint distributions of absolute magnitude in the $r$ band, $M_{r}$ (as an indicator of inherent galaxy brightness) versus $g-r$ galaxy color. As before, blue-dashed and black-solid represent the unweighted distribution for all binary hosts and the distribution weighted by the GW detectability of the constituent MBHB, respectively. The orange-dotted line shows the distribution in the color-magnitude plane for all galaxies, not just binary hosts. Immediately obvious is that galaxies harboring MBHBs that are detectable in GWs by PTAs tend to be brighter and redder than typical binary hosts and typical galaxies overall. This also implies that these galaxies are quenched ellipticals. We observe that the distribution of binary host properties overlaps with those of all galaxies and the hosts of GW-detectable binaries (as one would expect), but that the compounding requirements of a galaxy hosting a binary that is also detectable in GWs by PTAs leads to virtually negligible overlap in the black-solid and orange-dotted distributions. It is worth noting that the difference between unweighted binary-hosts and all galaxies is partially due to selection cuts: recall that in Illustris MBH are seeded into halos above a certain mass, which will correlate with brightness.
@@
The red dash-dotted line in Figure 7 indicates the decision boundary for a linear support vector machine (SVM) classifier [e.g. 113] between these latter two distributions, achieving virtually perfect classification (i.e., the area under the receiver operator characteristic curve was $1.0$) on $20\%$ of sampled systems that were held out from training. [^6] We note that the SVM was trained on equal numbers of sampled systems from each distribution, such that the classification was based on the distribution of properties and not rates. By contrast, a similar SVM trained to discriminate between binary hosts and all galaxies achieves a true positive rate of $\sim 65\%$ with a false positive rate of $0.1\%$, due to the larger (visible) overlap in those distributions. A full systematic investigation of the optimal classification scheme for which galaxy observables can aid in discriminating the hosts of detectable MBHBs is beyond the scope of this paper, but our proof-of-principle analysis here does demonstrate the feasibility of doing so with color–magnitude and other galaxy information.